# Complexity Matters: Dynamics of Feature Learning in the Presence of Spurious Correlations

### v1

@@ 4 Theoretical Explanation
This implies that the core feature will continue to increase but at a slower convergence rate which depends on the correlation strength $\lambda$. Therefore, if the spurious feature is simpler, then the core feature gradient reduces to the lower value earlier, leading to even slower convergence rate. This is consistent with our empirical observations. Note that our lemma can only show that core gradient will be lower, but does not give us the actual value of the gradient at this point.
@@
Another empirical observation is the persistence of spurious features despite the core feature being learned. Here we present a justification for this in an idealized setup. We consider a neural network that can be decomposed into two sub-networks, $h_{s}$ and $h_{c}$, such that $h(x)=h_{s}(x)+h_{c}(x)$ where $h_{s}(x)=\sum_{i\in S}a_{i}\sigma(w_{i}^{\top}x_{s}),h_{c}(x)=\sum_{i\in C}b_{i}\sigma(v_{i}^{\top}x_{c})$. The $S,C$ represent the index set of spurious and core neurons respectively. Note that we base this assumption on the empirical finding we have in (R3). We further make the assumption that both the spurious feature and core feature have been learned sufficiently well (as observed in our experiments). In particular we assume the following:
@@
Consequently, the population loss of the model on $\mathcal{D}_{\lambda}$ is given by the equation:
@@
The loss function is convex when $\lambda$ is within the range $[0,1]$ with respect to $\gamma_{s}$. The optimal point occurs when $\gamma_{c}\to\infty$, at which point $\gamma_{s}\to 0$. However, in practice, due to bounded iterations of optimization, this ideal scenario is unattainable. Instead, $\gamma_{c}$ will inevitably be less than some constant. In such cases, a Bayes optimal model that considers both features will have a positive value for $\gamma_{s}$ given that the spurious feature has been learned in the early stage. fig. 6 illustrates the optimal value of $\gamma_{s}$ for varying $\lambda$ when $\gamma_{c}$ is fixed at a constant. [30] studied the dynamics of core and spurious features in a similar form under a linear 2D setting and showed that the convergence rate of spurious weight to 0 is exponentially slow.
@@
Figure 6: Optimal values of spurious margin under different $\lambda$ with varying confidence core margin. The left plot shows how spurious weights might changes with varying $\lambda$ and a fixed $\alpha$, which implies retraining a model with more balanced data points would yield most significant improvement when $\lambda$ is high and core feature is learned poorly. The right plot shows weights on learned spurious feature might not decrease significantly after it is learned (even when the core feature is learned sufficiently well). Note that under the cross-entropy loss, the margin represents the confidence level of a model, hence $\textbf{confidence}=\phi(\gamma)\leq 1$. For $\gamma_{c}=15$, we have $\textbf{confidence}>0.99999$, which is likely to be an upper bound for most iterative optimization methods. The plots are generated by fixing $\lambda$ and $\gamma_{c}$ at various values and optimizing $\gamma_{s}$ with respect to $L_{\mathcal{D}_{\lambda}}$.
@@
The left plot in fig. 6 offers insights into the efficacy of LLR (or data balancing retraining in general) when both the spurious and core features can be adequately learned. Incorporating minority group data points effectively diminishes the convergence point of the spurious margin by reducing $\lambda$, subsequently decreasing the weights of the spurious subnetwork, as evidenced by our findings. Additionally, the plot illustrates that LLR yields more pronounced improvements when the confidence in the core feature is low, as observed in (R4) during the early stages of learning. This observation aligns with the findings in [25], where the addition of a small number of minority examples results in the most significant improvement. The right plot demonstrates that $\gamma_{s}$ does not decrease significantly within the feasible range of $\gamma_{c}$. It is noteworthy that the values of $\gamma_{s}$ at $\gamma_{c}=0$ are equal to the value calculated in lemma 2, thereby offering an estimate of the weight dynamics within the spurious network after lemma 2.
@@ 5 Conclusion
On the theoretical side, studying the end-to-end dynamics with spurious features is a technically interesting problem, and may require new tools. Our analysis currently assumes certain well-motivated conditions to bypass some parts of the unknown underlying feature learning process, and we believe that rigorously proving them would be challenging. Notably, showing that the spurious and core networks remain disjoint seems particularly non-trivial to show. A key finding of our analysis is the persistence of spurious feature weights, which tend to converge to significant values when $\lambda$ is high even if the model has learned the core feature confidently. Our investigation reveals that retraining the model with more balanced dataset effectively reduces the weights on the spurious network. While we refer to the observed correlation as “spurious" in this work, it’s important to acknowledge that there are scenarios where we seek to learn a diverse array of features.
Acknowledgement.
We thank Ben L. Edelman, Vaishnavh Nagarajan, and the anonymous reviewers of the M3L workshop at NeurIPS for their helpful suggestions to improve the paper.
@@ A.3 Properties of Boolean spurious distribution
[…]
@@ C.2.1 Complexity
Figure 14: Learning dynamics of Domino dataset. With core image as CiFar Truck-AutoMobile and spurious feature as the title of each subplot. $\lambda=0.95$ is fixed.
@@ C.2.2 Confounder Strength
(figs. 18 and 19) Surprisingly, our observations indicate that the pretrained model exhibits not only insensitivity to spurious features across a spectrum of complexities but also a remarkable resistance to higher $\lambda$ values. Additionally, when compared to the initialization with random weights, models with pretrained weights consistently maintain low spurious correlations throughout the training process.
@@ C.3.1 More neuron plots on a 2 layer NN
Figure 20: Dynamics of neurons on Parity task. The color of each lines is corresponding to the legend shown in fig. 3.
@@
Figure 21: Dynamics of neurons on Staircase task. The color of each lines is corresponding to the legend shown in fig. 3.
@@ C.3.3 Spurious/Core neurons on Vision Datasets
Dataset Core Spurious Core w.o Spurious Spurious w.o Sore
@@ D.1 Limitations of Existing Debiasing Algorithms
Figure 25: Core/spurious correlation and decoded correlation dynamics of different datasets. Dataset specification is outlined in fig. 24
@@ D.2 Discussion on spurious real dataset
Real-world datasets typically used to study spurious correlations, such as MultiNLI[47] and Civilcomment[7], often do not meet the properties of the boolean spurious datasets where core feature and spurious feature are disentangled and realizable. In these datasets, spurious attributes are intricately intertwined at the word level, such that the removal of a negative word can significantly alter a sentence’s semantic meaning. Furthermore, it has been reported that the labels in these datasets are not entirely clear-cut and may be inherently ambiguous, posing challenges for semantic labeling even for human annotators, thereby violating the realizability condition. We observe that the training dynamics in such real-world datasets are diversified and not fully understood, as indicated by [15, 16] (refer to fig. 25).

### v2

@@ 4 Theoretical Explanation
This implies that the core feature will continue to increase but at a slower convergence rate which depends on the correlation strength $\lambda$. Therefore, if the spurious feature is simpler, then the core feature gradient reduces to the lower value earlier, leading to even slower convergence rate. This is consistent with our empirical observations.
@@
Another empirical observation is the persistence of spurious features despite the core feature being learned. Here we present a justification for this in an idealized setup. We consider a neural network that can be decomposed into two sub-networks, $h_{s}$ and $h_{c}$, such that $h(x)=h_{s}(x)+h_{c}(x)$ where $h_{s}(x)=\sum_{i\in S}a_{i}\sigma(w_{i}^{\top}x_{s}),h_{c}(x)=\sum_{i\in C}b_{i}\sigma(v_{i}^{\top}x_{c})$. The $S,C$ represent the index set of spurious and core neurons respectively. Note that we base this assumption on the empirical finding we have in (R3).
We further make the assumption that the spurious feature is learned and being Bayes optimal throughout the later stage of training (as observed in our experiments) while the core feature is being learned and the model gives homogeneous response to $x_{c}$. In particular we assume the following:
@@
The reason we require that the model gives homogesnous response to $x_{c}$ is due to the property of cross entropy loss which has caused the slow down to be non-linear for different samples if response to $x_{c}$ is not the same.
We first show that spurious neurons are "dead" as they do not learn core feature in the later stage.
Lemma 4 (informal).
If $\sum_{i\in C}|w_{i}|<|w_{j}|$ for all $j\in S$. Then the gradient on core coordinates of the spurious neuron will be 0.
The lemma would also suggest that learned spurious neurons occupy part of the neural network capacity and if the network is of small size, the core feature may not be learned at all.
On the other hand, the population loss of the model on $\mathcal{D}_{\lambda}$ is given by the equation:
@@
The loss function is convex when $\lambda$ is within the range $[0,1]$ with respect to $\gamma_{s}$. The optimal point occurs when $\gamma_{c}\to\infty$, at which point $\gamma_{s}\to 0$. However, in practice, due to bounded iterations of optimization, this ideal scenario is unattainable. Instead, $\gamma_{c}$ will inevitably be less than some constant. In such cases, a Bayes optimal model that considers both features will have a positive value for $\gamma_{s}$ given that the spurious feature has been learned in the early stage. We conduct numerical experiment (see fig. 7) to illustrates the optimal value of $\gamma_{s}$ for varying $\lambda$ and $\gamma_{c}$ and the slow down ratio of core feature gradient. The slow down ratio can be formulated as $\frac{2\lambda(1-\phi(\gamma_{s}^{*}+\gamma_{c}))}{1-\phi(\gamma_{c})}$ where the numerator is gradient toward core feature under spurious distribution and the denominator is the same when spurious and core feature is uncorrelated or $\lambda=0.5$. We use $\gamma_{s}^{*}$ to denote the optimal value of spurious margin under a fixed $\gamma_{c}$ and $\lambda$.
@@
Figure 7: Optimal values of spurious margin under different $\lambda$ with varying confidence core margin. The left plot shows how spurious weights changes with varying $\lambda$ and confidence on core feature. The plot implies retraining a model with more balanced data points would yield most significant improvement when $\lambda$ is high and core feature is learned poorly. The right plot shows how core feature learning is slowed down by the learned spurious feature. Note that under the cross-entropy loss, the margin represents the confidence level of a model, hence $\textbf{confidence}=\phi(\gamma)\leq 1$. For $\gamma_{c}=15$, we have $\textbf{confidence}>0.99999$, which is likely to be an empirical upper bound for most iterative optimization methods. The plots are generated by fixing $\lambda$ and $\gamma_{c}$ at various values and optimizing $\gamma_{s}$ with respect to $L_{\mathcal{D}_{\lambda}}$.
@@
The left plot in fig. 7 offers insights into the efficacy of LLR (or data balancing retraining in general) when both the spurious and core features can be adequately learned. Incorporating minority group data points effectively diminishes the convergence point of the spurious margin by reducing $\lambda$, subsequently decreasing the weights of the spurious subnetwork, as evidenced by our findings. Additionally, the plot illustrates that LLR yields more pronounced improvements when the confidence in the core feature is low, as observed in (R4) during the early stages of learning. This observation aligns with the findings in [28], where the addition of a small number of minority examples results in the most significant improvement. The right plot demonstrates that the slow down is less significant in the later stage of training although $\lambda$ plays a key role in determine the empirical convergence point. It is noteworthy that the values of $\gamma_{s}$ at $\gamma_{c}=0$ are equal to the value calculated in lemma 2, thereby offering an estimate of the weight dynamics within the spurious network after lemma 2.
@@ 5 Conclusion
On the theoretical side, studying the end-to-end dynamics with spurious features is a technically interesting problem, and may require new tools. Our analysis currently assumes certain well-motivated conditions to bypass some parts of the unknown underlying feature learning process, and we believe that rigorously proving them would be challenging. Notably, showing that the spurious and core networks remain disjoint seems particularly non-trivial to show. A key finding of our analysis is the persistence of spurious feature weights, which tend to converge to significant values when $\lambda$ is high even if the model has learned the core feature confidently. Our investigation reveals that retraining the model with more balanced dataset effectively reduces the weights on the spurious network. While we refer to the observed correlation of another feature as "spurious" in this context, it’s essential to recognize that there are instances where correlated features can be benign and offer valuable insights into the core feature. In such cases, we may aim to learn both the spurious and core features.
Acknowledgements
We thank Ben L. Edelman, Vaishnavh Nagarajan, and the anonymous reviewers of the M3L workshop at NeurIPS and ICML for their helpful suggestions to improve the paper.
@@ A.3 Properties of Boolean spurious distribution
When $f_{c}(x_{c})$ is unbiased, $x_{s}$ is marginally uniformly distributed on $\mathcal{D}_{\text{same}},\mathcal{D}_{\text{diff}},\mathcal{D}_{\lambda}$. And when $f_{s}(x_{s})$ is unbiased $x_{c}$ is marginally uniformly distributed on $\mathcal{D}_{\text{same}},\mathcal{D}_{\text{diff}},\mathcal{D}_{\lambda}$
@@ B.2 Dynamics after spurious feature has been learned
Lemma B.5.
Let $I_{c}$ denote core coordinates indexes and $I_{s}$ denote spurious coordinates indexes. Suppose a neural network can be decomposed into $h_{s}(x_{s})+h_{c}(x_{c})$. Further, suppose the core network gives the same prediction $\gamma_{c}f_{c}(x_{c})$ on all $x_{c}$ i.e $h_{c}(x_{c})=\beta,\forall x_{c}$. If $\sum_{i\in I_{c}}|w_{i}|<|w_{j}|$ for all $j\in I_{s}$. Then the gradient on core coordinates of the spurious neuron will be 0.
Proof.
We have for any weight on a core coordinate $c\in I_{c}$, $\nabla_{w_{c}}L_{\mathcal{D}_{\mathsf{\lambda}}}(m(x))=-\lambda\nabla_{w_{c}}E_{\mathcal{D}_{\mathsf{same}}}[\log(\phi(f_{s}(x_{s})m(x))]-(1-\lambda)\nabla_{w_{c}}E_{\mathcal{D}_{\mathsf{diff}}}[\log(\phi(-f_{s}(x_{s})m(x))]$. We will show $\nabla_{w_{c}}E_{\mathcal{D}_{\mathsf{same}}}[\log(\phi(f_{s}(x_{s})m(x))]=0$ and the same can be shown in a similar manner for $\mathcal{D}_{\mathsf{diff}}$.
$\nabla_{w_{c}}E_{\mathcal{D}_{\mathsf{same}}}[\log(\phi(f_{s}(x_{s})m(x))]$
$=$ $E_{\mathcal{D}_{\mathsf{same}}}[(1-\phi(f_{s}(x_{s})h_{s}(x_{s})+\gamma_{c})f_{s}(x_{s})a\sigma^{\prime}(w^{\top}x)x_{c}]$
$=$ $2aE_{\mathcal{D}_{\mathsf{unif}}}[(1-\phi(f_{s}(x_{s})h_{s}(x_{s})+\gamma_{c})f_{s}(x_{s})\frac{f_{s}(x_{s})f_{c}(x_{c})+1}{2}x_{c}]$
$=$ $aE_{\mathcal{D}_{\mathsf{unif}}}[(1-\phi(f_{s}(x_{s})h_{s}(x_{s})+\gamma_{c})(f_{s}(x_{s})+f_{c}(x_{c}))\sigma^{\prime}(w_{s}^{\top}x_{s})x_{c}]$
$=$ $aE_{\mathcal{D}_{\mathsf{unif}}}[(1-\phi(f_{s}(x_{s})h_{s}(x_{s})+\gamma_{c})\chi_{I_{c}\setminus\{c\}}\sigma^{\prime}(w_{s}^{\top}x_{s})]+aE_{\mathcal{D}_{\mathsf{unif}}}[(1-\phi(f_{s}(x_{s})h_{s}(x_{s})+\gamma_{c})f_{s}(x_{s})\sigma^{\prime}(w_{s}^{\top}x_{s})x_{c}]$
$=$ $0+0$
@@ C.2.1 Complexity
Figure 15: Learning dynamics of Domino dataset. The spurious dataset become harder from left to right. The subplot title shows the spurious task. mnist01: Classification task of handwritten digits images of 0 and 1 taken from the MNIST dataset. mnist79: Classification task of handwritten digits images of 7 and 9 taken from the MNIST dataset. fashion dress coat: Classification task of dress and coat images taken from the FashionMnist dataset. The core task is classification of Truck and AutoMobile images taken from the CiFar dataset. The plots shows similar to the staircase dataset, harder spurious feature has less influence of the end performance of the model. Note confounder strength $\lambda=0.95$ is fixed.
@@ C.2.2 Confounder Strength
(figs. 19 and 20) Surprisingly, our observations indicate that the pretrained model exhibits not only insensitivity to spurious features across a spectrum of complexities but also a remarkable resistance to higher $\lambda$ values. Additionally, when compared to the initialization with random weights, models with pretrained weights consistently maintain low spurious correlations throughout the training process.
@@ C.3.1 More neuron plots on a 2 layer NN
Figure 21: Dynamics of neurons on Parity task.
@@
Figure 22: Dynamics of neurons on Staircase task.
@@ C.3.3 Spurious/Core neurons on Vision Datasets
Dataset Core Spurious Core w.o Spurious Spurious w.o core
@@ D.1 Limitations of Existing Debiasing Algorithms
Figure 26: Training dynamics of different spurious datasets with default pretrained weights: Dataset specification is outlined in fig. 25
Figure 27: Training dynamics of different spurious datasets with random weights: Dataset specification is outlined in fig. 25
@@ D.2 Discussion on spurious real dataset
Real-world datasets typically used to study spurious correlations, such as MultiNLI[49] and Civilcomment[7], often do not meet the properties of the boolean spurious datasets where core feature and spurious feature are disentangled and realizable. In these datasets, spurious attributes are intricately intertwined at the word level, such that the removal of a negative word can significantly alter a sentence’s semantic meaning. Furthermore, it has been reported that the labels in these datasets are not entirely clear-cut and may be inherently ambiguous, posing challenges for semantic labeling even for human annotators, thereby violating the realizability condition. We observe that the training dynamics in such real-world datasets are diversified and not fully understood, as indicated by [16, 17] (refer to fig. 26, fig. 27). It lacks justifications whether these datasets are suitable to be used in studying spurious correlation.